1) When Uranium-235 (atomic number 92/mass number 235) is struck by a slow neutron it is split into two smaller elements and 3 neutrons (atomic number 0/mass number 1). If one of the elements it splits into is Strontium (Sr) (atomic number 38/mass number 90), what is the other element it splits into? Must include atomic number and mass number.  

2) Uranium-238 eventually decays into Lead-206. The half-life of Uranium-238 is 4.5 x 109 or 4.5 billion years. A rock found on earth contains 50% Uranium-238 and 50% Lead-206. How old is that rock?

3) Nickel (atomic number 28, mass number 62) when it undergoes beta decay, it becomes what new element? Give both the name of the new element, new atomic number, and new mass number.

4) Francium (atomic number 87 and mass number 223) when it undergoes alpha radiation, it becomes what new element? Give both the name of the new element, new atomic number, and new mass number.
Answer #1

These are the question than can be solved by using laws of radioactivity which state That whenever a radioactive unstable nucleus decay through emitting an alpha particle having 4 units of mass and 2 units of charged the resultant nucleus has a mass decrease by 4 units and charge decrease by 2 units. Similarly, when an unstable nucleus is decay by emitting a beta particle than the resultant nucleus has a charge unit rise by 1units and mass remains the same. During this radioactive decay process, the mass and charge of the parent unstable nucleus remain conserved.

1. Uranium has a mass no. 235 and charge no.92 if it split to two-parent atoms by emitting 3 neutrons having mass 1 unit and 0 charge unit than it converts to Sr having mass no.90 and charge no. 38 by adding all the mass of Sr and neutron we have 90+3 = 93 so the Uranium must have to split into an element having 235-93= 142 mass no. and 92-42= 50 atomic no. So, the element having 142 mass no. and 50 atomic no. is Stannous (see mass no. can be more or less than the natural Stannous).

2. In this case, the age of the rock will similar to its half-life periods because about 50% of Uranium 235 has been decay. Half-life period is the time needed to decay 50% of a radioactive nucleus. So, the age of the rock will 4.5 billion years.

3. The decay of Nickel-62 having atomic no. 28 will follow the law of radioactivity and through beta decay its mass no. remains the same but atomic no. rise by one unit so the new element form will have mass no. 62 and atomic no. 29 and the new element will Copper.

4. Similarly, Francium 223 having atomic no. 87 when undergoes alpha decay the new element form has a mass no. 223-4 = 219 and atomic no 87-2= 85 the new element will Astatine.